# thanks: Corresponding Author

### v1

@@ Abstract
Hadronic molecules $\mathfrak{M}=J/\psi B_{c}^{+}$ and $\widetilde{\mathfrak{M}}=\eta_{c}B_{c}^{\ast+}$ are investigated in the framework of QCD sum rule method. These particles with spin-parities $J^{\mathrm{P}}=1^{+}$ have the quark contents $cc\overline{c}\overline{b}$. We compute their masses and current couplings and find that they are numerically very close to each other. Because it is difficult to distinguish reliably the molecules $J/\psi B_{c}^{+}$ and $J/\psi B_{c}^{+}$, we treat them as identical structures, and consider in details the state $\mathfrak{M}$. Our prediction $m=(9740\pm 70)~\mathrm{MeV}$ for its mass means that $\mathfrak{M}$ easily decays to pairs of ordinary mesons through strong interactions. There are two mechanisms responsible for transformations of $\mathfrak{M}$ to conventional mesons. The fall-apart mechanism generates the dominant decay channels $\mathfrak{M}\to J/\psi B_{c}^{+}$ and $\mathfrak{M}\to\eta_{c}B_{c}^{\ast+}$. Annihilation of $\overline{c}c$ quarks triggers subdominant processes with various final-state $B$ and $D$ mesons: Six of such channels are explored in this work. The partial widths of all decays are computed using the three-point sum rule approach. The width $\Gamma[\mathfrak{M}]=(121\pm 17)~\mathrm{MeV}$ of the hadronic axial-vector molecule $\mathfrak{M}$, as well as its mass are valuable for running and future experiments.
@@ I Introduction
Fully heavy hadronic molecules, i.e., molecules composed of heavy mesons and containing only $b$ and $c$ quarks are relatively new states attracted interests due to recent LHCb-CMS-ATLAS discoveries of $X$ structures. These presumably exotic $cc\overline{c}\overline{c}$ mesons were investigated by employing different approaches. The diquark-antidiquark and hadronic molecule pictures are mostly utilized models to describe properties of these states. Relevant problems were also addressed in our articles Refs. Agaev:2023ruu ; Agaev:2023rpj , in which we computed parameters of the molecules $\eta_{c}\eta_{c}$, $\chi_{c0}\chi_{c0}$, and $\chi_{c1}\chi_{c1}$ and interpreted some of them as candidates to $X$ structures.
@@
The heavy charm-bottom molecules built of equal number of $b$ and $c$ quarks were examined in different papers Liu:2023gla ; Liu:2024pio ; Wang:2023bek ; Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw . The coupled-channel unitary approach was applied in Ref. Liu:2023gla to study states $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$. The parameters of such molecules with $J^{\mathrm{PC}}=0^{++}$, $1^{++}$, and $2^{++}$ were calculated in our works Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw . We performed corresponding analyses using the QCD sum rule (SR) method, which allows one to evaluate parameters of ordinary and exotic hadrons Shifman:1978bx ; Shifman:1978by ; Albuquerque:2018jkn ; Agaev:2020zad . This approach permits us not only to find parameters of hadrons, for example the mass $m$, but also to estimate ambiguities $\pm\delta m$ emerged after applied approximations.
@@
The fully heavy asymmetric molecules with contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ were explored in the literature Liu:2024pio ; Agaev:2025did ; Agaev:2025wyf , as well. In Refs. Agaev:2025did ; Agaev:2025wyf the scalar $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$, and axial-vector $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{c}B_{c}^{+}]$ structures were studied using the SR method, where we computed the masses and decay widths of these structures. The molecules with a prevalence of $b$-quarks are close to corresponding two-meson thresholds provided one uses for analysis the central value $m$. In this scenario, both $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{AV}}$ are resonant states and dissociate to constituent mesons. In the second scenario $m-\delta m$ the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{AV}}$ reside below two-meson thresholds and form bound structures. Nevertheless, even in this case these states can decay to ordinary mesons. Such processes occur due to annihilation of $b\overline{b}$ quarks in $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{AV}}$ into light $q\overline{q}$ and $s\overline{s}$ pairs followed by creations of ordinary mesons Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara . Stated differently, there are two mechanisms for transformation of four-quark hadronic molecules to ordinary particles. Both of them were used to estimate total decay widths of the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{AV}}$.
The scalar molecule $\mathcal{M}_{\mathrm{c}}$ is above the $\eta_{c}B_{c}^{+}$ threshold and stays there at $m-\delta m$. Hence, it does not form a bound state. In present work, we study the axial-vector counterparts of $\mathcal{M}_{\mathrm{c}}$, i.e., consider the molecules $\mathfrak{M}=J/\psi B_{c}^{+}$ and $\widetilde{\mathfrak{M}}=\eta_{c}B_{c}^{\ast+}$ built of $cc\overline{c}\overline{b}$ quarks. It is evident that these quarks may be grouped into two different colorless components-ordinary heavy mesons. Thus, $\mathfrak{M}$ is a system of the vector charmonium $J/\psi$ and pseudoscalar $B_{c}^{+}$ meson, whereas $\widetilde{\mathfrak{M}}$ contains the pseudoscalar charmonium $\eta_{c}$ and vector meson $B_{c}^{\ast+}$.
We are going to compute the masses and decay widths of $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$ in the framework of QCD SR method. It turns out that whilst $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$ have different internal organizations their masses and current couplings are numerically very close each other. As a result, we consider $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$ as molecules indistinguishable within the context of SR method and examine features of $\mathfrak{M}$.
@@
In this work, we address also various decay channels of the molecule $\mathfrak{M}$. It is unstable against strong decays to ordinary meson pairs through two mechanisms discussed just above. For the molecule $\mathfrak{M}$ the processes $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are its dominant decay modes. The second mechanism gives rise to subdominant channels. In the case under consideration, we concentrate on the decays $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$, $B^{\ast 0}D^{+}$, $B^{+}D^{\ast 0}$$B^{0}D^{\ast+}$, $B_{s}^{\ast 0}D_{s}^{+}$, and $B_{s}^{0}D_{s}^{\ast+}$ which belong to this category of processes.
@@
This article is organized in the following way:, we compute The masses and current couplings of the hadronic molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are calculated in Sec. II. In Sec. III we consider the decay channels $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ and calculate their partial widths. The subdominant modes of $\mathfrak{M}$ are explored in sections IV and V. In section V we present our prediction for the full decay width of $J/\psi B_{c}^{+}$. The last part VI of this paper is reserved for discussion and contains also our concluding remarks.
@@ II Masses and current couplings of the molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p)$ $=$ $\frac{\langle 0|I_{\mu}|\mathfrak{M}(p,\epsilon)\rangle\langle\mathfrak{M}(p,\epsilon)|I_{\nu}^{\dagger}|0\rangle}{m^{2}-p^{2}}$ (4)
$+\cdots,$
@@
[…]
$\left.\times S_{c}^{ba^{\prime}}(x)\gamma_{\nu}S_{c}^{a^{\prime}a}(-x)\right]-\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{aa^{\prime}}(x)\gamma_{\nu}S_{c}^{a^{\prime}a}(-x)\right]$
$\left.\times\mathrm{Tr}\left[\gamma_{5}S_{c}^{bb^{\prime}}(x)\gamma_{5}S_{b}^{b^{\prime}b}(-x)\right]\right\},$ (7)
@@
To start numerical computations we fix the gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle=(0.012\pm 0.004)~\mathrm{GeV}^{4}$ and masses $m_{c}=(1.2730\pm 0.0046)~\mathrm{GeV}$ and $m_{b}=(4.183\pm 0.007)~\mathrm{GeV}$ of the quarks are universal quantities Shifman:1978bx ; Shifman:1978by ; PDG:2024 . The parameters $M^{2}$ and $s_{0}$ depend on a considering process and have to satisfy usual constraints of SR studies. The prevalence of the pole contribution ($\mathrm{PC}$) in extracted quantities is one of such important restrictions. Therefore, we choose $M^{2}$ and $s_{0}$ in such a way that to ensure $\mathrm{PC}\geq 0.5$ . The next constraint is convergence of the operator product expansion. We calculate $\Pi(M^{2},s_{0})$ by taking into account dimension-$4$ nonperturbative terms, therefore require fulfillment of $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$ is sufficient for convergence of $\mathrm{OPE}$. Additionally, physical parameters of $\mathfrak{M}$ have to be stable upon variations of $M^{2}$ and $s_{0}$.
@@
Figure 1: $\mathrm{PC}$ as a function of the Borel parameter $M^{2}$ at fixed $s_{0}$. The star shows the point $M^{2}=9~\mathrm{GeV}^{2}$ and $s_{0}=110~\mathrm{GeV}^{2}$.
@@
$m$ $=$ $(9740\pm 70)~\mathrm{MeV},$
$\Lambda$ $=$ $(5.19\pm 0.49)~\times 10^{-1}\text{ }\mathrm{GeV}^{5}.$ (12)
@@
$\widetilde{m}$ $=$ $(9737\pm 71)~\mathrm{MeV},$
$\widetilde{\Lambda}$ $=$ $(5.23\pm 0.50)~\times 10^{-1}\text{ }\mathrm{GeV}^{5}.$ (13)
It is clear that these parameters are very close to ones written down in Eq. (12): There are wide overlapping regions for $m$, $\Lambda$ and $\widetilde{m}$, $\widetilde{\Lambda}$. Because theoretical errors do not allow us to distinguish the states $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$, we treat them as identical structures and, in what follows, explore the molecule $\mathfrak{M}$.
@@ III.1 $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$
$\Pi_{\mu\nu}^{1}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I^{B_{c}}(y)$ (14)
$\times I_{\mu}^{J/\psi}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$
@@
The correlator $\Pi_{\mu\nu}^{1}(p,p^{\prime})$ written down using masses and decay constants (current coupling) of the particles $\mathfrak{M}$, $J/\psi$, and $B_{c}^{+}$ forms the phenomenological side $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ of SR. After dissecting the contribution of the ground-state particles, it has the form
$\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|I^{B_{c}}|B_{c}^{+}(p^{\prime})\rangle}{p^{\prime 2}-m_{B_{c}}^{2}}\frac{\langle 0|I_{\mu}^{J/\psi}|J/\psi(q,\varepsilon)\rangle}{q^{2}-m_{J/\psi}^{2}}$
$\times\langle J/\psi(q,\varepsilon)B_{c}^{+}(p^{\prime})|\mathfrak{M}(p,\epsilon)\rangle\frac{\langle\mathfrak{M}(p,\epsilon)|I_{\nu}^{\dagger}|0\rangle}{p^{2}-m^{2}}$
$+\cdots,$ (16)
@@
$\langle 0|I_{\mu}^{J/\psi}|J/\psi(q,\varepsilon)\rangle$ $=$ $f_{J/\psi}m_{J/\psi}\varepsilon_{\mu},$
$\langle 0|I^{B_{c}}|B_{c}^{+}(p^{\prime})\rangle$ $=$ $\frac{f_{B_{c}}m_{B_{c}}^{2}}{m_{b}+m_{c}},$ (17)
@@
$\langle J/\psi(q,\varepsilon)B_{c}^{+}(p^{\prime})|\mathfrak{M}(p,\epsilon)\rangle$ $=$ $g_{1}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})\right$. (18)
$\left.-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$
@@
$\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})=\frac{g_{1}(q^{2})\Lambda f_{B_{c}}m_{B_{c}}^{2}f_{J/\psi}m_{J/\psi}}{(m_{b}+m_{c})\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{B_{c}}^{2})}$
$\times\frac{1}{(q^{2}-m_{J/\psi}^{2})}\left[\frac{m^{2}+m_{B_{c}}^{2}-q^{2}}{2}g_{\mu\nu}-\frac{m^{2}}{m_{J/\psi}^{2}}p_{\mu}^{\prime}p_{\nu}^{\prime}\right$.
$\left.+\frac{m^{2}-m_{J/\psi}^{2}}{m_{J/\psi}^{2}}p_{\mu}^{\prime}p_{\nu}-\frac{m^{2}+m_{B_{c}}^{2}-q^{2}}{2m_{J/\psi}^{2}}(p_{\mu}p_{\nu}+p_{\mu}p_{\nu}^{\prime})\right]$
$+\cdots.$ (19)
@@
[…]
$\left.\times\gamma_{\nu}S_{c}^{aj}(x)\right]\mathrm{Tr}\left[\gamma_{5}S_{c}^{ib}(y-x)\gamma_{5}S_{b}^{bi}(x-y)\right]$
$\left.-\mathrm{Tr}\left[\gamma_{5}S_{c}^{ia}(y-x)\gamma_{\nu}S_{c}^{aj}(x)\gamma_{\mu}S_{c}^{jb}(-x)\gamma_{5}S_{b}^{bi}(x-y)\right]\right\}.$
@@
$g_{1}(q^{2})=\frac{2(m_{b}+m_{c})}{\Lambda f_{B_{c}}m_{B_{c}}^{2}f_{J/\psi}m_{J/\psi}}\frac{q^{2}-m_{J/\psi}^{2}}{m^{2}+m_{B_{c}}^{2}-q^{2}}$
$\times e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}}^{2}/M_{2}^{2}}\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (21)

### v2

@@ Abstract
Hadronic molecules $\mathfrak{M}=J/\psi B_{c}^{+}$ and $\widetilde{\mathfrak{\ M}}=\eta_{c}B_{c}^{\ast+}$ are investigated in the framework of QCD sum rule method. These particles with spin-parities $J^{\mathrm{P}}=1^{+}$ have the quark contents $cc\overline{c}\overline{b}$. We compute their masses and current couplings and find that they are numerically very close to each other coinciding within accuracy of the sum rule method. Therefore, we concentrate on the molecule $J/\psi B_{c}^{+}$ and explore features of this state in a detailed form. Our prediction $m=(9740\pm 70)~\mathrm{MeV}$ for its mass means that $\mathfrak{M}$ easily decays to pairs of ordinary mesons through strong interactions. There are two mechanisms responsible for transformations of $\mathfrak{M}$ to conventional mesons. The fall-apart mechanism generates the dominant decay channels $\mathfrak{M}\to J/\psi B_{c}^{+}$ and $\mathfrak{M}\to\eta_{c}B_{c}^{\ast+}$. Annihilation of $\overline{c}c$ quarks triggers subdominant processes with various final-state $B$ and $D$ mesons: Six of such channels are considered in this work. The partial widths all of decays are computed using the three-point sum rule approach. The width $\Gamma[\mathfrak{M}]=(121\pm 17)~\mathrm{MeV}$ of the hadronic axial-vector molecule $\mathfrak{M}$, as well as its mass provide valuable information for running and future experiments.
@@ I Introduction
Fully heavy hadronic molecules, i.e., molecules composed of heavy mesons and containing only $b$ and $c$ quarks are relatively new states attracted interests due to recent LHCb-CMS-ATLAS discoveries of $X$ structures LHCb:2020bwg ; ATLAS:2023bft ; CMS:2023owd . These presumably exotic $cc\overline{c}\overline{c}$ mesons were investigated by employing different approaches. The diquark-antidiquark and hadronic molecule pictures are mostly utilized models to describe properties of these states. Relevant problems were also addressed in our articles Refs. Agaev:2023ruu ; Agaev:2023rpj , in which we computed parameters of the molecules $\eta_{c}\eta_{c}$, $\chi_{c0}\chi_{c0}$, and $\chi_{c1}\chi_{c1}$ and interpreted some of them as candidates to $X$ structures.
@@
The heavy charm-bottom molecules built of equal number of $b$ and $c$ quarks were examined in different papers Liu:2023gla ; Liu:2024pio ; Wang:2023bek ; Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw . The coupled-channel unitary approach was applied in Ref. Liu:2023gla to study states $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$. The parameters of such molecules with $J^{\mathrm{PC}}=0^{++}$, $1^{++}$, and $2^{++}$ were calculated in our works Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw . We performed corresponding analyses using the QCD sum rule (SR) method, which allows one to evaluate parameters of ordinary and exotic hadrons Shifman:1978bx ; Shifman:1978by ; Albuquerque:2018jkn ; Agaev:2020zad . In this approach we find the mass $m\pm\delta m$ of a particle with the central value $m$ and ambiguities $\pm\delta m$ emerged during these computations. In other words, the region $m-\delta m$, $m+\delta m]$ can be used to fix the mass of a particle of interest.
@@
The fully heavy asymmetric molecules with contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ were explored in the literature Liu:2024pio ; Agaev:2025did ; Agaev:2025wyf as well. In Refs. Agaev:2025did ; Agaev:2025wyf the scalar $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$, and axial-vector $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{b}B_{c}^{\ast+}$] structures were studied using the SR method, where we computed the masses and decay widths of these structures. The molecules with a prevalence of $b$-quarks are close to corresponding two-meson thresholds provided one uses for analysis the central value $m$. In this scenario, both the scalar $\mathcal{M}_{\mathrm{b}}$ and axial-vector $\mathcal{M}_{\mathrm{AV}}$ molecules are resonant states and dissociate to constituent mesons. In the lower limit of the masses, i.e., in a situation when $m-\delta m$, the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{AV}}$ reside below two-meson thresholds and form bound structures. This means that they cannot decay to mesons which are their ingredients, or decay to particles which contain all four initial quarks. Nevertheless, $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{AV}}$ transform to ordinary mesons through strong processes which occur due to annihilation of $b\overline{b}$ quarks into light $q\overline{q}$ and $s\overline{s}$ pairs followed by creations of ordinary mesons Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara . Thus, there are two mechanisms for transformation of four-quark hadronic molecules to ordinary particles. In our previous works Agaev:2025did ; Agaev:2025wyf we used both of these mechanisms to evaluate total widths of the hadronic molecules $\eta_{b}B_{c}^{-}$ and $\Upsilon B_{c}^{-}$.
In the case of the scalar molecule $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$ this scenario changes qualitatively. It turns out that the molecule $\mathcal{M}_{\mathrm{c}}$ dissociates to its components $\eta_{c}+B_{c}^{+}$ even in the lower value of its mass. Hence, it does not form a bound state in a sense emphasized above.
In present work, we study the axial-vector counterparts of $\mathcal{M}_{\mathrm{c}}$, i.e., consider the molecules $\mathfrak{M}=J/\psi B_{c}^{+}$ and $\widetilde{\mathfrak{M}}=\eta_{c}B_{c}^{\ast+}$ built of $cc\overline{c}\overline{b}$ quarks. It is evident that these quarks may be grouped into two different colorless components-ordinary heavy mesons. Thus, $\mathfrak{M}$ is a system of the vector charmonium $J/\psi$ and pseudoscalar $B_{c}^{+}$ meson, whereas $\widetilde{\mathfrak{M}}$ contains the pseudoscalar charmonium $\eta_{c}$ and vector meson $B_{c}^{\ast+}$. We are going to compute the masses and decay widths of $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$ in the framework of QCD SR method. We find that whilst $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$ have different internal organizations and are different molecules, their masses and current couplings are numerically very close each other. They have identical quark contents and quantum numbers, as a result, their decay patterns do not differ from each other as well. Because the SR method leads to predictions with some uncertainties, we cannot clearly distinguish these structures. Therefore, in this work we concentrate on properties of the molecule $\mathfrak{M}$.
@@
To this end, apart from the mass and current coupling, we address also various decay channels of the molecule $\mathfrak{M}$. It is strong-interaction unstable state and decays to ordinary meson pairs through two mechanisms discussed just above. The processes $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are dominant decay modes of the molecule $\mathfrak{M}$. The second mechanism gives rise to subdominant channels $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$, $B^{\ast 0}D^{+}$, $B^{+}D^{\ast 0}$$B^{0}D^{\ast+}$, $B_{s}^{\ast 0}D_{s}^{+}$, and $B_{s}^{0}D_{s}^{\ast+}$ which belong to this category of processes.
@@
This article is organized in the following way: The masses and current couplings of the hadronic molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ are calculated in Sec. II. In Sec. III we consider the decay channels $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$ and calculate their partial widths. The subdominant modes of $\mathfrak{M}$ are explored in Secs. IV and V. In section V we present our prediction for the full decay width of the molecule $J/\psi B_{c}^{+}$. The last part, VI, of this paper is reserved for discussion and contains also our concluding remarks.
@@ II Masses and current couplings of the molecules $J/\psi B_{c}^{+}$ and $\eta_{c}B_{c}^{\ast+}$
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p)=\frac{\langle 0|I_{\mu}|\mathfrak{M}(p,\epsilon)\rangle\langle\mathfrak{M}(p,\epsilon)|I_{\nu}^{\dagger}|0\rangle}{m^{2}-p^{2}}+\cdots,$ (4)
@@
$\Pi_{\mu\nu}^{\mathrm{OPE}}(p)$ $=$ $i\int d^{4}xe^{ipx}\mathrm{Tr}\left\{\left[\gamma_{\mu}S_{c}^{ab^{\prime}}(x)\gamma_{5}S_{b}^{b^{\prime}b}(-x)\gamma_{5}S_{c}^{ba^{\prime}}(x)\gamma_{\nu}S_{c}^{a^{\prime}a}(-x)\right]\right$. (7)
$\left.-\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{aa^{\prime}}(x)\gamma_{\nu}S_{c}^{a^{\prime}a}(-x)\right]\mathrm{Tr}\left[\gamma_{5}S_{c}^{bb^{\prime}}(x)\gamma_{5}S_{b}^{b^{\prime}b}(-x)\right]\right\},$
@@
To start numerical computations we fix the gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle=(0.012\pm 0.004)~\mathrm{GeV}^{4}$ and masses $m_{c}=(1.2730\pm 0.0046)~\mathrm{GeV}$ and $m_{b}=(4.183\pm 0.007)~\mathrm{GeV}$ of the quarks which are universal quantities Shifman:1978bx ; Shifman:1978by ; PDG:2024 . The parameters $M^{2}$ and $s_{0}$ depend on a considering process and have to satisfy usual constraints of SR studies. The prevalence of the pole contribution ($\mathrm{PC}$) in extracted quantities is one of such important restrictions. Therefore, we choose $M^{2}$ and $s_{0}$ in such a way that to ensure $\mathrm{PC}\geq 0.5$ . The next constraint is convergence of the operator product expansion. We calculate $\Pi(M^{2},s_{0})$ by taking into account dimension-$4$ nonperturbative terms, therefore require fulfillment of $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$ which is sufficient for convergence of $\mathrm{OPE}$. Additionally, physical parameters of $\mathfrak{M}$ have to be stable upon variations of $M^{2}$ and $s_{0}$.
@@
Figure 1: $\mathrm{PC}$ as a function of the Borel parameter $M^{2}$ at fixed $s_{0}$. The star shows the point $M^{2}=9~\mathrm{GeV}^{2},\ s_{0}=110~\mathrm{GeV}^{2}$.
@@
$m=(9740\pm 70)~\mathrm{MeV},\ \ \Lambda=(5.19\pm 0.49)~\times 10^{-1}\text{ }\mathrm{GeV}^{5}.$ (12)
@@
$\widetilde{m}=(9737\pm 71)~\mathrm{MeV},\ \ \widetilde{\Lambda}=(5.23\pm 0.50)~\times 10^{-1}\text{ }\mathrm{GeV}^{5}.$ (13)
It is clear that these parameters are very close to ones written down in Eq. (12): There are wide overlapping regions for $m$, $\Lambda$ and $\widetilde{m}$, $\widetilde{\Lambda}$. Because theoretical errors do not allow us to distinguish clearly the states $\mathfrak{M}$ and $\widetilde{\mathfrak{M}}$, in what follows, we explore the molecule $\mathfrak{M}$.
@@ III.1 $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$
$\Pi_{\mu\nu}^{1}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I^{B_{c}}(y)I_{\mu}^{J/\psi}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$ (14)
@@
The correlator $\Pi_{\mu\nu}^{1}(p,p^{\prime})$ written down using masses and decay constants (current coupling) of the particles $\mathfrak{M}$, $J/\psi$, and $B_{c}^{+}$ forms the phenomenological side $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ of SR. After dissecting the contribution of the ground-state particles in the factorization approximation Ioffe:1982qb ; Colangelo:2000dp ; Agaev:2022iha , it has the form
$\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ $=$ $\frac{\langle 0|I^{B_{c}}|B_{c}^{+}(p^{\prime})\rangle}{p^{\prime 2}-m_{B_{c}}^{2}}\frac{\langle 0|I_{\mu}^{J/\psi}|J/\psi(q,\varepsilon)\rangle}{q^{2}-m_{J/\psi}^{2}}\langle J/\psi(q,\varepsilon)B_{c}^{+}(p^{\prime})|\mathfrak{M}(p,\epsilon)\rangle$ (16)
$\times\frac{\langle\mathfrak{M}(p,\epsilon)|I_{\nu}^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
$\langle 0|I_{\mu}^{J/\psi}|J/\psi(q,\varepsilon)\rangle=f_{J/\psi}m_{J/\psi}\varepsilon_{\mu},\ \ \langle 0|I^{B_{c}}|B_{c}^{+}(p^{\prime})\rangle=\frac{f_{B_{c}}m_{B_{c}}^{2}}{m_{b}+m_{c}},$ (17)
@@
$\langle J/\psi(q,\varepsilon)B_{c}^{+}(p^{\prime})|\mathfrak{M}(p,\epsilon)\rangle=g_{1}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$ (18)
@@
$\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})=\frac{g_{1}(q^{2})\Lambda f_{B_{c}}m_{B_{c}}^{2}f_{J/\psi}m_{J/\psi}}{(m_{b}+m_{c})\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{B_{c}}^{2})}\frac{1}{(q^{2}-m_{J/\psi}^{2})}\left[\frac{m^{2}+m_{B_{c}}^{2}-q^{2}}{2}g_{\mu\nu}\right$.
$\left.-\frac{m^{2}}{m_{J/\psi}^{2}}p_{\mu}^{\prime}p_{\nu}^{\prime}+\frac{m^{2}-m_{J/\psi}^{2}}{m_{J/\psi}^{2}}p_{\mu}^{\prime}p_{\nu}-\frac{m^{2}+m_{B_{c}}^{2}-q^{2}}{2m_{J/\psi}^{2}}(p_{\mu}p_{\nu}+p_{\mu}p_{\nu}^{\prime})\right]+\cdots.$ (19)
@@
$\Pi_{\mu\nu}^{1\mathrm{OPE}}(p,p^{\prime})=-\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\left\{\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ja}(-x)\gamma_{\nu}S_{c}^{aj}(x)\right]\mathrm{Tr}\left[\gamma_{5}S_{c}^{ib}(y-x)\gamma_{5}S_{b}^{bi}(x-y)\right]\right$.
$\left.-\mathrm{Tr}\left[\gamma_{5}S_{c}^{ia}(y-x)\gamma_{\nu}S_{c}^{aj}(x)\gamma_{\mu}S_{c}^{jb}(-x)\gamma_{5}S_{b}^{bi}(x-y)\right]\right\}.$ (20)
@@
$g_{1}(q^{2})=\frac{2(m_{b}+m_{c})}{\Lambda f_{B_{c}}m_{B_{c}}^{2}f_{J/\psi}m_{J/\psi}}\frac{q^{2}-m_{J/\psi}^{2}}{m^{2}+m_{B_{c}}^{2}-q^{2}}e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}}^{2}/M_{2}^{2}}\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (21)